OpenStudy (anonymous):

Suppose a gas tank is in the shape of a cylinder. The tank is 20 feet long and 8 feet in diameter. Find the amount of metal used to construct the tank. 384.5 ft3 1028 ft3 602.2 ft3 1570.4 ft3

OpenStudy (rajee_sam):

it should be ft^2. The amount of metal is the surface area actually

OpenStudy (anonymous):

What?

OpenStudy (rajee_sam):

the units in your answers they are not correct

OpenStudy (anonymous):

Its on the problem that the teacher wrote out, so I dont think so.

OpenStudy (rajee_sam):

you ask your teacher tomorrow she will change it. P.S> Teachers can be wrong sometimes. they are humans too

OpenStudy (anonymous):

Heheh...he's right. "Cubic feet" is a volumetric measurement. The metal used to enclose an area is a surface, and should be "square feet". It may be just a typo on your teachers part. Hold on and I'll check.

OpenStudy (rajee_sam):

\[Surface Area of Cylinder = (2 \times \pi \times r ^{2}) + (2 \times \pi \times r \times h)\]

OpenStudy (rajee_sam):

use the data given and plug in the values in the formula and find your answer.

OpenStudy (anonymous):

I got 602.8 How come my choice of answers only include 602.2?

OpenStudy (anonymous):

What are you using for pi?

OpenStudy (anonymous):

3.14?

OpenStudy (anonymous):

Okay, hold on... :-)

OpenStudy (anonymous):

Isnt the radius the diameter cut in half?

OpenStudy (anonymous):

I got 602.88 using 3.14 And yes, radius is half the diameter.

OpenStudy (anonymous):

So would it be answer C?

OpenStudy (anonymous):

Yeah, I'd go with C

OpenStudy (anonymous):

BTW, I answered your last question in the other thread, too, just in case you thought I forgot about it... :-)

OpenStudy (anonymous):

Okay, thank you:)